# Linearization in magnetoelasticity

### v1

@@ 4. Proof of $\Gamma$-convergence
By Proposition 3.1, there exist $u\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$ and $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence, $u_{\varepsilon}\rightharpoonup u$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $m_{\varepsilon}\circ y_{\varepsilon}\,\det\nabla y_{\varepsilon}\to m$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $1\leq r<\infty$, $m_{\varepsilon}\circ y_{\varepsilon}\to m$ in $L^{1}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$, and $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then, by Proposition 2.1 we have that
$\int_{\Omega}|\nabla m|^{2}{\,\rm d}x+\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}x\leq\liminf_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}x+\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}x.$
@@
[…]
Thus, $\chi_{E_{\varepsilon}}\to 1$ in measure and $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u$ in $L^{2}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Indeed, for $21\leq r<$ 2 we may estimate
@@
Therefore, $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}=\nabla u_{\varepsilon}-\chi_{\Omega\setminus E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u\text{ weakly in }L^{r}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Since the weak limit in $L^{r}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ is unique, we get that $\chi_{E_{\varepsilon}}\nabla u\rightharpoonup\nabla u$ weakly in $L^{2}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$
@@
$\geq\frac{1}{{2}}\int_{E_{\varepsilon}}\mathbb{C}(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})):(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})){\,\rm d}x$
@@
(4.9) $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}\text{ in }L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),$
@@
The injectivity of $y_{\varepsilon}$ and the sign condition $\det\nabla y_{\varepsilon}>0$ follow from Theorem 2.2 ([8, Theorem 5.5-1]), since $\varepsilon\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla u_{0}\|_{\infty}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq c(\Omega)$ for small enough $\varepsilon$.
Let us now notice that $y_{\varepsilon}$ is bi-Lipschitz for small enough $\varepsilon$. Indeed, there exist constants $L$ and $l>0$ (that may depend on $u_{0}$ but not on $x$ and $\varepsilon$), such that for small enough $\varepsilon$ it holds that $\sup_{x\in\Omega}|\nabla y_{\varepsilon}|=\sup_{x\in\Omega}|I+\varepsilon\nabla u_{0}(x)|=L_{\varepsilon}<L$ with $L_{\varepsilon}\to 1$ if $\varepsilon\to 0$. Moreover, $\inf_{x\in\Omega}\det\nabla y_{\varepsilon}=\inf_{x\in\Omega}\det(I+\varepsilon\nabla u_{0})=l_{\varepsilon}>l$ and $l_{\varepsilon}\to 1$ if $\varepsilon\to 0$. By the inverse function theorem $y_{\varepsilon}^{-1}$ is differentiable and $\nabla y_{\varepsilon}^{-1}(y_{\varepsilon}(x))=(\nabla y_{\varepsilon}(x))^{-1}\approx I-\varepsilon\nabla u_{0}(x)$. Thus, $m_{\varepsilon}$ defined in (4.5) can be explicitly written as
@@
Therefore, taking the squares, integrating over $\Omega_{\delta}$, and letting $\varepsilon\to 0$ we infer that $m_{\varepsilon}\to m_{0}$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$, for every $\delta>0$. Arguing as in the proof of Proposition 3.1, it is not difficult to see that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Moreover, by (4.11) we can estimate the exchange energy as
$\limsup_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}(y)|^{2}{\,\rm d}y$ $\leq\limsup_{\varepsilon\to 0}\int_{\Omega}\frac{L_{\varepsilon}^{3d-2}}{l_{\varepsilon}^{4}}|\nabla m_{0}(x)|^{2}{\,\rm d}x+C\frac{\|\nabla\det\nabla y_{\varepsilon}\|^{2}_{L^{\infty}}}{l_{\varepsilon}^{2}}=\int_{\Omega}|\nabla m_{0}(x)|^{2}{\,\rm d}x.$
@@
$\limsup_{\varepsilon\to 0}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})$ $=\limsup_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\int_{\Omega}W(I+\varepsilon\nabla u_{0},m_{0}){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}z$
@@
We conclude this section with the proof of convergence of minimizers of $\mathcal{F}_{\varepsilon}$ to minimizers of $\mathcal{F}_{0}$, namely, Theorems 1.2 and 1.3. We start with the coercivity of $\mathcal{F}_{\varepsilon}$.
@@
$\|u_{\varepsilon}\|_{L^{2}(\Omega;\mathbb{R}^{d})}^{2}\leq C\bigg{(}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}\,{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
It is standard to show that $\mathcal{F}_{0}$ has a minimizer $(u_{0},m_{0})\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ on $\mathcal{A}_{0}^{w}$. It is also straightforward to check that $\inf\mathcal{F}_{\varepsilon_{k}}(u,m)$ is bounded w.r.t. $\varepsilon$. Since the functionals $\mathscr{L}$ and $\mathscr{M}$ do not depend on $\varepsilon$, from $\Gamma$-convergence of $\mathcal{G}_{\varepsilon}$ we deduce that $s_{\varepsilon_{k}}\to s_{0}$ and thus $\mathcal{F}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{F}_{0}(u_{0},m_{0})$. Propositions 4.1 and 3.1 imply that $u_{\varepsilon_{k}}\rightharpoonup u_{0}$ weakly in $W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon_{k}(\Omega)}}m_{\varepsilon_{k}}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon_{k}(\Omega)}}\nabla m_{\varepsilon_{k}}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\mathcal{G}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{G}(u_{0},m_{0})$. ∎
@@
The convergence of the sequence $u_{\varepsilon}$ can be improved to a strong convergence in $W^{1,2}(\Omega;\mathbb{R}^{d})$ arguing as in [16, Subsection 7.2] (see also [6, Subsection 3.2]).
Acknowledgements
The work of S.A. was funded by the Austrian Science Fund (FWF) through the projects P 35359-N and ESP 65, by the Italian Ministry of Education and Research through the PRIN 2022 project “Variational Analysis of Complex Systems in Material Science, Physics and Biology” No. 2022HKBF5C, and by the University of Naples Federico II through the FRA project “Regularity and Singularity in Analysis, PDEs, and Applied Sciences”. S.A. also acknowledges the hospitality of the University of Vienna and the TU Wien, where part of this work has been done. M.K. was supported by the GAČR-FWF project 21-06569K and by the Ministry of Education, Youth and Sport of the Czech Republic through the bilateral project 8J22AT017. A.M. was supported by the European Unions Horizon 2020 research and innovation programme under the Marie Składowska-Curie grant agreement No. 847693.
Finally, S.A. and A.M. acknowledge the kind hospitality of the ESI Institute (Vienna) during the workshop “New perspective on Shape and Topology Optimization”.

### v2

@@ 4. Proof of $\Gamma$-convergence
By Proposition 3.1, there exist $u\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$ and $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence, $u_{\varepsilon}\rightharpoonup u$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $m_{\varepsilon}\circ y_{\varepsilon}\,\det\nabla y_{\varepsilon}\to m$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $1\leq r<\infty$, $m_{\varepsilon}\circ y_{\varepsilon}\to m$ in $L^{1}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$, and $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then, by Proposition 2.1 we have that
$\int_{\Omega}|\nabla m|^{2}{\,\rm d}x+\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}x\leq\liminf_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}z+\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}z.$
@@
$\mathcal{L}^{d}(\Omega\setminus E_{\varepsilon})=\mathcal{L}^{d}\big{(}\left\{x\in\Omega:|\nabla u_{\varepsilon}(x)|\geq\varepsilon^{-\alpha}\right\}\big{)}\leq\frac{1}{\varepsilon^{-\alpha}}\int_{\Omega\setminus E_{\varepsilon}}|\nabla u_{\varepsilon}(x)|{\rm d}x\leq\varepsilon^{\alpha}|\Omega|^{1/2}\|\nabla u_{\varepsilon}\|_{L^{2}}\leq C\varepsilon^{\alpha}.$
Thus, $\chi_{E_{\varepsilon}}\to 1$ in measure and $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u$ in $L^{2}(\Omega;\mathbb{R}^{d\times d})$. Indeed, for $21\leq r<$ 2 we may estimate
@@
Therefore, $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}=\nabla u_{\varepsilon}-\chi_{\Omega\setminus E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u\text{ weakly in }L^{r}(\Omega;\mathbb{R}^{d\times d})$. Since the weak limit is unique, we get that $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u$ weakly in $L^{2}(\Omega;\mathbb{R}^{d\times d})$.
@@
$=\frac{1}{{2}}\int_{E_{\varepsilon}}\mathbb{C}(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})):(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})){\,\rm d}x$
@@
(4.9) $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}\text{ in }L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d}),$
@@
The injectivity of $y_{\varepsilon}$ and the sign condition $\det\nabla y_{\varepsilon}>0$ follow from Theorem 2.2 ([15, Theorem 5.5-1]), since $\varepsilon\|\nabla u_{0}\|_{\infty}\leq c(\Omega)$ for small enough $\varepsilon$. Moreover, $y_{\varepsilon}$ is bi-Lipschitz for small enough $\varepsilon$. Indeed, there exist constants $0<l<L<+\infty$ (that may depend on $u_{0}$ but not on $x$ and $\varepsilon$), such that for $\varepsilon>0$ small enough it holds that $\sup_{x\in\Omega}|\nabla y_{\varepsilon}|=\sup_{x\in\Omega}|I+\varepsilon\nabla u_{0}(x)|\leq 1+\varepsilon\sup_{x\in\Omega}|\nabla u_{0}(x)|=:L_{\varepsilon}<L$ with $L_{\varepsilon}\to 1$ if $\varepsilon\to 0$. At the same time, $\inf_{x\in\Omega}\det\nabla y_{\varepsilon}=\inf_{x\in\Omega}\det(I+\varepsilon\nabla u_{0})=1+\varepsilon\inf_{x\in\Omega}\operatorname{tr}\nabla u_{0}(x)+o(\varepsilon)=:l_{\varepsilon}>l>0$ and $l_{\varepsilon}\to 1$ if $\varepsilon\to 0$. By the inverse function theorem $y_{\varepsilon}^{-1}$ is differentiable and $\nabla y_{\varepsilon}^{-1}(y_{\varepsilon}(x))=(\nabla y_{\varepsilon}(x))^{-1}\approx I-\varepsilon\nabla u_{0}(x)$. Thus, $m_{\varepsilon}$ defined in (4.5) can be explicitly written as
@@
Therefore, taking the squares, integrating over $\Omega_{\delta}$, and letting $\varepsilon\to 0$ we infer that $m_{\varepsilon}\to m_{0}$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$, for every $\delta>0$. Arguing as in the proof of Proposition 3.1, it is not difficult to see that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d\times d})$. Moreover, by (4.11) we can estimate the exchange energy as
$\limsup_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}(y)|^{2}{\,\rm d}z$ $\leq\limsup_{\varepsilon\to 0}\int_{\Omega}\frac{L_{\varepsilon}^{3d-2}}{l_{\varepsilon}^{4}}|\nabla m_{0}(x)|^{2}{\,\rm d}x+C\frac{\|\nabla\det\nabla y_{\varepsilon}\|^{2}_{L^{\infty}}}{l_{\varepsilon}^{2}}=\int_{\Omega}|\nabla m_{0}(x)|^{2}{\,\rm d}x.$
@@
$\limsup_{\varepsilon\to 0}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})$ $=\limsup_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\int_{\Omega}W(I+\varepsilon\nabla u_{0},m_{0}){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}z+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}z$
@@
We conclude this section with the proof of convergence of minimizers of $\mathcal{F}_{\varepsilon}$ to minimizers of $\mathcal{F}_{0}$, namely, Theorems 1.3 and 1.4. We start with the coercivity of $\mathcal{F}_{\varepsilon}$.
@@
$\|u_{\varepsilon}\|_{L^{2}(\Omega;\mathbb{R}^{d})}^{2}\leq C\bigg{(}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})+\int_{\Gamma}|w|^{2}\,{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
@@
It is standard to show that $\mathcal{F}_{0}$ has a minimizer $(u_{0},m_{0})\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ on $\mathcal{A}_{0}^{w}$. It is also straightforward to check that $\inf\mathcal{F}_{\varepsilon_{k}}(u,m)$ is bounded with respect to $\varepsilon$. Since the functionals $\mathscr{L}$ and $\mathscr{M}$ do not depend on $\varepsilon$, from $\Gamma$-convergence of $\mathcal{G}_{\varepsilon}$ we deduce that $s_{\varepsilon_{k}}\to s_{0}$ and thus $\mathcal{F}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{F}_{0}(u_{0},m_{0})$. Propositions 4.1 and 3.1 imply that $u_{\varepsilon_{k}}\rightharpoonup u_{0}$ weakly in $W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon_{k}(\Omega)}}m_{\varepsilon_{k}}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon_{k}(\Omega)}}\nabla m_{\varepsilon_{k}}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\mathcal{G}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{G}(u_{0},m_{0})$. ∎
@@
The convergence of the sequence $u_{\varepsilon}$ can be improved to a strong convergence in $W^{1,2}(\Omega;\mathbb{R}^{d})$ arguing as in [27, Subsection 7.2] (see also [13, Subsection 3.2]). ∎
Acknowledgments
The work of S.A. was funded by the Austrian Science Fund (FWF) through the projects P 35359-N and ESP 65, by the Italian Ministry of Education and Research through the PRIN 2022 project “Variational Analysis of Complex Systems in Material Science, Physics and Biology” No. 2022HKBF5C, and by the University of Naples Federico II through the FRA project “Regularity and Singularity in Analysis, PDEs, and Applied Sciences”. S.A. also acknowledges the hospitality of the University of Vienna and the TU Wien, where part of this work has been done. M.K. was supported by the GAČR-FWF project 21-06569K, GAČR project 23-04766S, and by the Ministry of Education, Youth and Sport of the Czech Republic through the bilateral project 8J22AT017. A.M. was supported by the European Unions Horizon 2020 research and innovation programme under the Marie Składowska-Curie grant agreement No. 847693. S.A. and A.M. acknowledge the kind hospitality of the ESI Institute (Vienna) during the workshop “New perspective on Shape and Topology Optimization”.
Finally, the authors would like to express their gratitude to the anonymous reviewers for careful reading and their comments and suggestions that helped to improve the manuscript.